Help please

Help please
Help please
Answer #1

1) acceleration due to gravity on a planet, g = G*M/R^2

here G is universal grvatational constant
M is mass of the planet

R is the radius of the planet


so, M = g*R^2/G

= 3.2*(2.35*10^6)^2/(6.67*10^-11)

= 2.65*10^23 kg

mass of the person, m = Weigh on earth/g_earth

= 384/9.8

= 39.2 kg

so, weight of the person the given planet, w = m*g_planet

= 39.2*3.2

= 125.4 N

2)
here

mass of ice, m_ice = ?

mass of water, m_water = 741 grams = 0.741 kg

Apply

heat lost ice = heat gained by water

m_ice*C_ice*(17-0) + m_ice*Lf + m_ice*C_water*(12 - 0) = m_water*C_water*(70-12)

m_ice*(C_ice*17 + Lf + C_water*12) = m_water*C_water*58

m_ice = m_water*C_water*58/(C_ice*17 + Lf + C_water*12)

= 0.741*4190*58/(2000*17 + 334*10^3 + 4190*12)

= 0.43 kg or 430 grams